1+4=5
2+5=12
3+6=21
5+8=?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52777)

.
1*4 + 1 = 5,

2*5 + 2 = 12,

3*6 + 3 = 21,

5*8 + 4 = 44.


Answer. 44.


Answer by MathTherapy(10551)


1+4=5
2+5=12
3+6=21
5+8=?
Why was this reposted?
Wasn't the following answer for "Miscellaneous_Word_Problems/1087688: Miscellaneous_Word_Problems/1087688:" good enough for you?
Or, was that someone else who posted it earlier?